# SoK: Large Language Model Copyright Auditing via Fingerprinting

### v2

@@ Abstract
The broad capabilities and substantial resources required to train Large Language Models (LLMs) make them valuable intellectual property, yet they remain vulnerable to copyright infringement, such as unauthorized use and model theft. LLM fingerprinting, a non-intrusive technique that extracts and compares the distinctive features from LLMs to identify infringements, offers a promising solution to copyright auditing. However, its reliability remains uncertain due to the prevalence of diverse model modifications and the lack of standardized evaluation. In this SoK, we present the first comprehensive study of LLM fingerprinting. We introduce a unified framework and formal taxonomy that categorizes existing methods into white-box and black-box approaches, providing a structured overview of the state of the art. We further propose LeaFBench, the first systematic benchmark for evaluating LLM fingerprinting under realistic deployment scenarios. Built upon mainstream foundation models and comprising 149 distinct model instances, LeaFBench integrates 13 representative post-development techniques, spanning both parameter-altering methods ($e.g.$, fine-tuning, quantization) and parameter-independent mechanisms ($e.g.$, system prompts, RAG). Extensive experiments on LeaFBench reveal the strengths and weaknesses of existing methods, thereby outlining future research directions and critical open problems in this emerging field. The code is available at https://github.com/shaoshuo-ss/LeaFBench.
@@ 1 Introduction
With the rapid advancement of Large Language Models (LLMs), they have been extensively deployed across diverse domains, including content creation [57], language translation [16], and code generation [19]. Unlike traditional classification models, the development of LLMs is considerably more resource-intensive, demanding large-scale datasets and substantial computational power [32, 57]. However, despite their high value, LLMs are increasingly exposed to serious copyright infringement risks. For open-source models, adversaries may exploit them in unauthorized contexts ($e.g.$, prohibited commercial applications) or fine-tune the models and falsely claim independent authorship [75]. For closed-source models, adversaries can also extract model parameters or replicate functionality through model stealing attacks [6, 38]. These threats highlight an urgent need for effective mechanisms to protect and audit the copyright of LLMs in practice.
@@
LLM copyright auditing aims to determine whether a suspicious model is derived from a copyright-reserved source model. Existing approaches can be broadly categorized into two classes: LLM watermarking and fingerprinting. In general, LLM watermarking [^1] Some existing works embed “fingerprints” into LLMs by modifying their parameters [64, 69, 4]. In this paper, consistent with the classic narrow definitions of model watermarking and fingerprinting [68, 48, 33], we classify such methods as LLM watermarking. refers to the intrusive injection of a developer-specific identifier into the model [64, 29], which can later be extracted and validated to verify ownership. While effective in principle, watermarking requires direct modification of the model ($e.g.$, fine-tuning), which not only degrades performance but also introduces considerable computational overhead. More critically, watermarking cannot be retroactively applied to LLMs that have already been released, thereby restricting its applicability in real-world settings.
@@
To overcome the limitations of watermarking, LLM fingerprinting has recently emerged [44, 79, 76, 78, 17]. In general, fingerprinting involves non-intrusively extracting distinctive features ($e.g.$, parameter statistics or output characteristics) from the model. By comparing the fingerprints of different models, one can assess their similarity. A high degree of similarity often indicates copyright infringement. Unlike watermarking, fingerprinting does not require model modification and is thus also feasible for auditing released models. Notably, LLM fingerprinting has already seen practical application. For example, it was employed to reveal that a student group at Stanford had plagiarized the MiniCPM model [23]. More recently, several studies suggested that an open-source MoE model may have been developed based on Qwen-2.5 [75, 81], attracting widespread attention. These real-world cases highlight the practical potential of LLM fingerprinting for copyright protection and model auditing.
@@
Despite its promise, designing an effective and reliable LLM fingerprinting method is non-trivial. In contrast to fingerprinting traditional classification models [9, 7, 50], LLM fingerprinting faces two fundamental challenges: (1) The generation and sampling processes of LLMs are inherently non-deterministic, and such randomness may hinder the extraction of stable and consistent fingerprints. (2) Modern LLMs typically incorporate a wide range of techniques and components prior to deployment, including parameter-altering approaches ($e.g.$, fine-tuning and quantization [28]) as well as parameter-independent mechanisms ($e.g.$, system prompts [14], and retrieval-augmented generation (RAG) [31]). These factors significantly complicate the reliability of fingerprinting. Consequently, while fingerprinting has demonstrated potential in certain auditing scenarios, a critical question remains: Are existing LLM fingerprinting methods truly reliable in practice?
@@
In this SoK, we present the first comprehensive study of LLM copyright auditing via fingerprinting, aiming to provide a unified framework for analyzing both existing and future methods. We formalize LLM fingerprinting as a two-stage protocol: an $\rm{Extract}$ function that derives a fingerprint from a source model, and a $\rm{Verify}$ function that determines whether the fingerprint is present in a suspicious model. This principled approach allows us to construct a clear taxonomy that systematizes the current landscape. Specifically, we categorize methods by the auditor’s access level into white-box and black-box approaches. White-box methods are further classified into static, forward-pass, and backward-pass techniques, according to their feature source. Black-box methods, in contrast, are divided into untargeted and targeted fingerprinting, depending on whether they rely on specific query–response pairs. Leveraging this taxonomy, we systematically review representative and state-of-the-art (SOTA) methods, clarifying the connections and distinctions between different techniques.
@@
Beyond reviewing existing works, we address the lack of fair and standardized evaluation by introducing the Large Language Model Fingerprinting Benchmark (LeaFBench), the first systematic benchmark for evaluating LLM fingerprinting. LeaFBench comprises 149 carefully constructed LLM instances, derived from various mainstream base models. It further integrates 13 representative post-development techniques, ranging from parameter-altering methods to parameter-independent optimizations. This design enables fair and rigorous evaluation of fingerprinting methods under realistic conditions. Using LeaFBench, we conduct a comprehensive experimental study of 8 SOTA fingerprinting techniques. Our empirical analysis reveals important insights into the strengths and limitations of existing methods, highlights promising research directions, and uncovers critical open problems for advancing LLM copyright auditing in this rapidly evolving domain.
@@
  • Benchmark for Standardized Evaluation. We develop LeaFBench, the first systematic benchmark dedicated to LLM fingerprinting. LeaFBench provides a standardized toolbox to evaluate fingerprinting methods under a diverse range of representative and realistic model modifications.
  • Comprehensive Empirical Study. We conduct extensive experiments on 8 representative LLM fingerprinting methods using LeaFBench. Our analysis distills key insights into the effectiveness and limitations of current approaches, explores their practical applicability, and identifies critical open challenges and future research directions.
@@ 2.1 Background
Large Language Models (LLM). An LLM is a complex computational system, typically built on the Transformer architecture [61] and characterized by its massive parameter scale and auto-regressive generation paradigm [16]. Its development generally follows a multi-stage pipeline. It begins with pre-training, during which the model is trained on large-scale corpora to acquire general linguistic knowledge and syntactic patterns. After pre-training, developers apply additional techniques to adapt or refine the model for specific tasks or deployment environments. Such adaptations may involve directly modifying the model parameters through approaches like fine-tuning [22] and compression [28]. Alternatively, they may steer model behavior at inference time using parameter-independent mechanisms, such as system prompts [14] or retrieval-augmented generation (RAG) [31].
LLM Copyright Infringements. Developing LLMs requires large-scale datasets and massive computational resources, leading LLMs to valuable digital assets for their owner [32]. However, these valuable LLMs face significant copyright infringement risks in their application. For instance, an adversary could illegally utilize an open-source model in unauthorized areas or simply fine-tune it and claim it as their own [75]. For closed-source models that do not disclose their model parameters, this threat also has not been alleviated. Carlini et al. [6] demonstrate that it is possible for an adversary to extract precise information from production black-box LLMs. These real-world examples of illicit reuse and direct parameter theft underscore that such threats are not merely theoretical. Therefore, protecting and auditing the copyright of LLMs has become critically important in LLM applications [46].
@@ 2.2 Problem Formulation
$\mathcal{A}(M_{s},M_{o})=\{\begin{aligned} 1,&\kern 5.0ptM_{s}\text{ is derived from }M_{o}\\ 0,&\text{ otherwise}\end{aligned}.$ (1)
Notably, the method for identifying such “suspicious models” (and then auditing them) is generally not included in LLM copyright auditing [44, 82]. The identification process may rely on non-technical indicators, such as observing similarities in performance, functionality, or parameter count. This area of investigation is beyond the scope of our paper.
@@ 2.3 Threat Model
In this section, we introduce the threat model of LLM copyright auditing. There are two parties involved, including the adversary and the model owner (or auditor).
Assumptions of the Adversary. The adversary seeks to improperly appropriate a model (referred to as the source model) originally developed by the model owner by deploying it for unauthorized purposes or falsely claiming independent authorship. The model released or deployed by the potential adversary is referred to as the suspicious model. Once in possession of the source model after pre-training, the adversary has full control over it and may freely apply any build-time adaptations or runtime optimization techniques.
@@
  • Full control over the source model. The owner inherently possesses complete access to their own model $M_{o}$. This includes its architecture, parameters, and potentially the original training data. They can interact with and modify the model without any restrictions.
  • Access to the suspicious model. This access may be white-box, allowing inspection of the suspicious model’s internal parameters, or black-box, where interaction is limited to API queries and observing the corresponding outputs.
@@ 2.4 LLM Copyright Auditing Paradigms
To fulfill LLM copyright auditing, two primary technical paradigms have emerged in the literature: LLM watermarking [48, 64, 65] and LLM fingerprinting [44, 72, 42]. These approaches differ fundamentally in whether they require modification of the model being protected.
LLM Watermarking. LLM watermarking [73] is an active approach that intrusively embeds a unique identifier, or watermark, into the source model $M_{o}$ before its release. This process involves a pair of algorithms: an embedding function and a verification function. The copyright can later be verified by checking for the presence of this embedded watermark in a suspicious model $M_{s}$. The formal definition is as follows.
@@
LLM watermarking involves embedding a specific signal or message ($i.e.$, $w$) into the source model, $M_{o}$. This creates a modified, watermarked model $M_{w}$, which is then released. To audit a suspicious model $M_{s}$, the model owner could verify whether the predefined signal $w$ can be detected within it.
Despite its effectiveness, LLM watermarking suffers from several practical limitations [35, 44]. (1) The embedding process inevitably modifies the model’s parameters, which can degrade its performance on primary tasks. (2) This process often requires substantial additional computational overhead, such as fine-tuning the model to embed the watermark. (3) Most critically, watermarking is not applicable to models that have already been trained and released, as it must be performed proactively. These limitations motivate the development of LLM fingerprinting as an alternative auditing solution [76].
LLM Fingerprinting. LLM fingerprinting is a passive approach that does not require modifying the model. Instead, it extracts a set of inherent yet distinctive characteristics that collectively serve as the model’s “fingerprint” [44, 79, 76]. The core idea is that any model derived from $M_{o}$ will preserve a statistically significant portion of this fingerprint.
@@
The primary advantage of LLM fingerprinting lies in its non-invasive nature. Since it does not alter the model, it introduces no performance degradation. The computational overhead is typically confined, which is often less intensive than the fine-tuning required for watermarking. Consequently, fingerprinting offers a more flexible and widely applicable paradigm for copyright auditing, especially for models that are already in the public domain [74, 44, 50]. As such, in this paper, we primarily focus on LLM fingerprinting techniques.
3 Toward the Principle of LLM Fingerprinting
In this section, we present a comprehensive and unified formulation, including the framework, design objectives, and taxonomy, of LLM fingerprinting. Notably, in this paper, we primarily focus on the fingerprinting for LLM, which is an emerging and significant topic in the generative AI era [32, 46].
3.1 Framework of LLM Fingerprinting
In this section, we introduce the fundamental framework of LLM fingerprinting. Generally, an LLM fingerprinting scheme can be defined by a formal, two-stage protocol. This protocol is built upon two fundamental functions: one is the extraction function $\text{Extract}(\cdot)$ for extracting a signature, and the other is the fingerprint verification function $\text{Verify}(\cdot)$ for verifying it against a suspicious model.

### v3

@@ Abstract
The broad capabilities and substantial resources required to train Large Language Models (LLMs) make them valuable intellectual property, yet they remain vulnerable to copyright infringement, such as unauthorized use and model theft. LLM fingerprinting, a non-intrusive technique that compares the distinctive features ($i.e.$, fingerprint) of LLMs to identify whether an LLM is derived from another, offers a promising solution to copyright auditing. However, its reliability remains uncertain due to the prevalence of diverse model modifications and the lack of standardized evaluation. In this SoK, we present the first comprehensive study of the emerging LLM fingerprinting. We introduce a unified framework and taxonomy that structures the field: white-box methods are classified based on their feature source as static, forward-pass, or backward-pass fingerprinting, while black-box methods are distinguished by their query strategy as either untargeted or targeted. Furthermore, we propose LeaFBench, the first systematic benchmark for evaluating LLM fingerprinting under realistic deployment scenarios. Built upon 7 mainstream foundation models and comprising 149 distinct model instances, LeaFBench integrates 13 representative post-development techniques, spanning both parameter-altering methods ($e.g.$, fine-tuning, quantization) and parameter-independent techniques ($e.g.$, system prompts, RAG). Extensive experiments on LeaFBench reveal the strengths and weaknesses of existing methods, thereby outlining future research directions and critical open problems in this emerging field. The code is available at https://github.com/shaoshuo-ss/LeaFBench.
@@ 1 Introduction
With the rapid advancement of Large Language Models (LLMs), they have been extensively deployed across diverse domains, including content creation [qwen2.5], language translation [grattafiori2024llama3], and code generation [he2025benchmarking]. Compared with traditional small models, the development of LLMs is considerably more resource-intensive, demanding large-scale datasets and computational resources [li2025rethinking, qwen2.5]. However, despite their high value, LLMs are increasingly exposed to serious copyright infringement risks. For open-source models, adversaries may exploit them in unauthorized contexts ($e.g.$, prohibited commercial applications) or fine-tune the models and falsely claim independent authorship [yoon2025intrinsic]. For closed-source models, adversaries can also extract model parameters or replicate functionality through model stealing attacks [carlini2024stealing, liu2025model]. These threats highlight an urgent need for effective mechanisms to protect and audit the copyright of LLMs in practice.
@@
LLM copyright auditing aims to determine whether a suspicious model is derived from a copyright-reserved source model. Existing techniques can be broadly categorized into two classes: LLM watermarking and fingerprinting [sun2023deep]. LLM watermarking [^1] Some existing works embed “fingerprints” into LLMs by modifying their parameters [xu2024instructional, yamabe2025mergeprint, cai2025utf]. In this paper, consistent with the classic narrow definitions of model watermarking and fingerprinting [xue2021intellectual, shao2025explanation, li2025move], we classify such methods as LLM watermarking. refers to the intrusive injection of a specific identifier ($i.e.$, watermark) into the model [xu2024instructional, krauss2024clearstamp], which can later be extracted and validated to verify ownership. While effective in principle, watermarking requires direct modification of the model ($e.g.$, fine-tuning), which may not only degrade model performance but also incur substantial computational overhead. More importantly, watermarking cannot be applied to protect LLMs that have already been released, substantially limiting its applicability in real-world auditing scenarios.
@@
To overcome the limitations of watermarking, LLM fingerprinting techniques have recently emerged [pasquini2025llmmap, zhang2025reef, zeng2024huref, gubri2024trap]. Fingerprinting involves non-intrusively extracting distinctive fingerprints ($e.g.$, parameter statistics or output characteristics) from the models. By comparing the fingerprints, one can assess the similarity of different LLMs, and a high degree of similarity often indicates copyright infringement. Unlike watermarking, fingerprinting does not require modifying models. As such, it has no impact on the model performance and is also feasible for auditing released models. Notably, LLM fingerprinting has already seen practical application. For example, it was employed to reveal that a student group at Stanford had plagiarized the MiniCPM model [hu2024minicpm]. More recently, several studies suggested that a famous open-source MoE model may have been developed based on Qwen-2.5 [yoon2025intrinsic, zhang2025matrix], attracting widespread attention. These real-world cases highlight the practical potential of LLM fingerprinting for copyright protection and auditing.
@@
Despite its promise, designing an effective and reliable LLM fingerprinting method is non-trivial. In contrast to fingerprinting classification models [chen2022teacher, chen2022copy, shao2025fit], LLM fingerprinting faces two fundamental challenges: (1) The generation processes of LLMs are inherently non-deterministic, and such randomness may hinder the extraction of stable and consistent fingerprints. (2) Modern LLMs typically incorporate various techniques and components before deployment, including parameter-altering approaches ($e.g.$, fine-tuning and quantization [jin2024comprehensive]) as well as parameter-independent mechanisms ($e.g.$, system prompts [giray2023prompt], and retrieval-augmented generation (RAG) [lewis2020retrieval]). These factors significantly complicate the reliability of fingerprinting. Consequently, while fingerprinting has demonstrated potential, a critical question remains: Are existing LLM fingerprinting methods truly reliable in practice?
@@
In this SoK, we present the first comprehensive study of the emerging LLM copyright auditing via fingerprinting, aiming to provide a unified framework for analyzing existing and future methods. We formalize LLM fingerprinting as a two-stage framework: an $\rm{Extract}$ function that extracts a fingerprint from a source model, and a $\rm{Verify}$ function that determines whether the fingerprint is present in a suspicious model. This allows us to construct a clear taxonomy that systematizes the current landscape. Specifically, we categorize methods by the auditor’s access into white-box and black-box approaches. White-box methods are further classified into static, forward-pass, and backward-pass techniques, according to their feature source. Black-box methods, in contrast, are divided into untargeted and targeted fingerprinting, depending on whether they rely on specific query–response pairs. Leveraging this taxonomy, we systematically review existing LLM fingerprinting methods, clarifying the connections and distinctions between different techniques.
@@
Beyond reviewing existing works, we address the lack of standardized evaluation by introducing the Large Language Model Fingerprinting Benchmark (LeaFBench), the first systematic benchmark for evaluating LLM fingerprinting. LeaFBench comprises 149 carefully constructed LLM instances, derived from various mainstream base models. It further integrates 13 representative post-development techniques, ranging from parameter-altering methods ($e.g.$, fine-tuning and model merging) to parameter-independent optimizations ($e.g.$, RAG and CoT). This design enables fair and comprehensive evaluation of fingerprinting methods under realistic conditions. Using LeaFBench, we conduct a comprehensive experimental study of 8 state-of-the-art (SOTA) fingerprinting techniques. Our empirical analysis reveals important insights into the strengths and limitations of existing methods, highlights promising research directions, and uncovers critical open problems for advancing LLM copyright auditing in this rapidly evolving domain.
@@
  • Benchmark for Standardized Evaluation. We develop LeaFBench, the first systematic benchmark dedicated to LLM fingerprinting. It provides a standardized toolbox to evaluate fingerprinting methods under a diverse range of representative and realistic model modifications.
  • Comprehensive Empirical Study. We conduct extensive experiments on 8 existing representative and SOTA LLM fingerprinting methods using our LeaFBench. Our analysis summarizes key insights into the effectiveness and limitations of existing solutions, explores their practical applicability, and identifies critical open challenges and future research directions.
@@ 2.1 Background
Large Language Models (LLM). An LLM is a complex computational system, typically built on the Transformer architecture and characterized by its massive parameter scale and auto-regressive generation paradigm [grattafiori2024llama3]. Its development generally follows a multi-stage pipeline. It begins with pre-training, during which the model is trained on large-scale corpora to acquire general linguistic knowledge and syntactic patterns. After pre-training, developers apply additional techniques to adapt or refine the model for specific tasks or deployment environments. Such adaptations may involve directly modifying the model parameters through approaches like fine-tuning [hu2022lora] and compression [jin2024comprehensive]. Alternatively, they may steer model behavior at inference time using parameter-independent mechanisms, such as utilizing different system prompts [giray2023prompt] or a retrieval-augmented generation (RAG) mechanism [lewis2020retrieval].
LLM Copyright Infringements. Developing LLMs requires large-scale datasets and massive computational resources, leading LLMs to valuable digital assets for their owner [li2025rethinking]. However, these valuable LLMs face significant copyright infringement risks in their application. An adversary could illegally utilize an open-source model in unauthorized areas or fine-tune it and claim the fine-tuned version as their self-developed model [yoon2025intrinsic]. For instance, a student group is reported to have plagiarized the MiniCPM model [hu2024minicpm]. For closed-source models that do not directly disclose their model parameters, this threat also has not been alleviated. Existing study [carlini2024stealing] also demonstrates that it is possible for an adversary to extract precise information from production black-box LLMs. These real-world examples of illicit reuse and direct parameter theft underscore that such threats are not merely theoretical. Therefore, protecting and auditing the copyright of LLMs has become critically important in LLM applications [ren2024sok].
@@ 2.2 Problem Formulation
$\mathcal{A}(M_{s},M_{o})=\Big\{\begin{aligned} 1,&\kern 5.0ptM_{s}\text{ is derived from }M_{o}\\ 0,&\text{ otherwise}\end{aligned}.$ (1)
In this context, a model $M_{s}$ is considered “derived from” $M_{o}$ if it inherits a substantial portion of its learned parameters, and consequently its functional behavior, from $M_{o}$. This relationship holds regardless of whether $M_{s}$ is a direct copy, an illicitly stolen model, or a subsequent model created by applying post-development modifications ($e.g.$, fine-tuning, quantization) to $M_{o}$. Collectively, the base model and its derivatives form a “model lineage”. The primary goal of LLM copyright auditing is to distinguish the independently developed models from true derivatives.
Notably, the method for identifying such “suspicious models” (and then auditing them) is generally not included in LLM copyright auditing [pasquini2025llmmap, zhang2024remark]. The identification may rely on non-technical indicators, such as observing similarities in performance, functionality, or parameter count. This area is beyond the scope of our SoK.
@@ 2.3 Threat Model
There are two parties involved in the threat model of LLM copyright auditing, including the adversary and the model owner (or auditor).
Assumptions of the Adversary. The adversary seeks to improperly appropriate a model (referred to as the source model) originally developed by another model owner by copying and deploying it for unauthorized purposes or falsely claiming independent authorship. The model released or deployed by the potential adversary is referred to as the suspicious model. Once in possession of the source model after pre-training, the adversary has full control over it and could freely apply different techniques to adapt or optimize the model. The techniques will be discussed in Section 6.1.
@@
  • Full control over the source model. The owner/auditor inherently possesses complete access to their own model $M_{o}$. This includes its architecture, parameters, and potentially the original training data. They can interact with and modify the model without any restrictions.
  • Access to the suspicious model. The owner/auditor has access to the suspicious model. This access may be white-box, allowing inspection of the suspicious model’s internal parameters, or black-box, where interaction is limited to API queries and observing the corresponding outputs.
@@ 2.4 LLM Copyright Auditing Paradigms
To fulfill LLM copyright auditing, two primary technical paradigms have emerged in the literature: LLM watermarking [shao2025explanation, xu2024instructional, xu2025mark] and LLM fingerprinting [pasquini2025llmmap, yang2024fingerprint, mcgovern2025your]. These approaches differ fundamentally in whether they require modification of the model being protected.
LLM Watermarking. LLM watermarking [shao2025explanation, xu2024instructional, xu2025mark] is an active approach that intrusively embeds a unique identifier ($i.e.$, watermark) into the source model $M_{o}$ before its release or deployment. This process involves a pair of algorithms: an embedding function and a verification function. The copyright can later be verified by checking for the presence of this embedded watermark in a suspicious model $M_{s}$. The formal definition is as follows.
@@
LLM watermarking involves embedding a specific signal or message $w$ into the source model, $M_{o}$. This creates a modified, watermarked model $M_{w}$, which is then released. To audit a suspicious model $M_{s}$, the model owner could verify whether the predefined signal $w$ can be extracted from it.
Despite its effectiveness, LLM watermarking suffers from several practical limitations [li2023protecting, pasquini2025llmmap]. (1) The embedding process inevitably modifies the model’s parameters, which can degrade its performance on primary tasks. (2) This process often requires substantial additional computational overhead, such as fine-tuning the model to embed the watermark. (3) Most critically, model watermarking is not applicable to models that have already been trained and released, as it must be performed beforehand. These limitations motivate the development of LLM fingerprinting as an alternative auditing solution [zeng2024huref].
LLM Fingerprinting. LLM fingerprinting is a post-hoc approach that does not require modifying the model. Instead, it extracts a set of inherent yet distinctive characteristics that collectively serve as the model’s “fingerprint” [pasquini2025llmmap, zhang2025reef, zeng2024huref]. The core idea is that any model derived from $M_{o}$ will preserve a statistically significant portion of this fingerprint. The definition is as follows.
@@
The primary advantage of LLM fingerprinting lies in its non-invasive nature. Since it does not alter the model, it introduces no performance degradation. The computational overhead is typically confined, which is often less intensive than the fine-tuning required for watermarking. Consequently, fingerprinting offers a more flexible and widely applicable paradigm for copyright auditing, especially for models that are already in the public domain [yao2023removalnet, pasquini2025llmmap, shao2025fit]. As such, in this paper, we primarily focus on LLM fingerprinting techniques.
3 Principle of LLM Fingerprinting
In this section, we present a comprehensive and unified formulation, including the framework, objectives, and taxonomy, of LLM fingerprinting.
3.1 LLM Fingerprinting Framework
Generally, an LLM fingerprinting scheme can be defined by a formal, two-stage protocol. This protocol is built upon two fundamental functions: one is the extraction function $\text{Extract}(\cdot)$ for extracting a signature, and the other is the fingerprint verification function $\text{Verify}(\cdot)$ for verifying it against a suspicious model.